OpenStudy (anonymous):

The dimensions of a rectangular prism are 3, 4, and 5. The volume of the prism is 7 12 24 60

OpenStudy (anonymous):

@Mertsj

OpenStudy (opcode):

Use the formula: \(a*b*c\) \(3*4*5 = 60\) So the answer is \(\huge 60\) Here is a site with the list of formulas for volume: http://www.math.com/tables/geometry/volumes.htm

OpenStudy (anonymous):

If the radius and height of a cylinder are multiplied by 3, the lateral area of the cylinder is multiplied by 6. 9. 12. 18.

OpenStudy (anonymous):

how about that one please help

OpenStudy (anonymous):

i postd it as a new question and tagged you. please

OpenStudy (opcode):

If the radius and the height are multiplied by 3, that means that \(k = 3\) Since are looking for the area, which is \(k^2\) \(k = 3\) Input the values: \(k^2 = 3^2 = 9\)

OpenStudy (anonymous):

A triangular pyramid has a volume of 25 ft³. If all its dimensions are doubled, the volume of the new pyramid will be _____ ft³. 50 150 200 400

OpenStudy (anonymous):

you are such a good helper! i appreciate it. and telling me how

OpenStudy (anonymous):

hello?

OpenStudy (opcode):

Hold on. A triangular pyramid has a volume of 25 ft³. If all its dimensions are doubled, the volume of the new pyramid will be _____ ft³. \(2^3 (25)= 200ft^3\)

OpenStudy (anonymous):

A cylinder has a radius of 4 cm and a height of 10 cm. The surface area of the cylinder is _____ cm². 40 80 112 224

OpenStudy (anonymous):

and pie sympbol is after each one of those answer options

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

i appreciate it

OpenStudy (opcode):

http://answers.yahoo.com/question/index?qid=20110503175553AAWTzcd Go to the link he explains it better. Anyway gotta go.